OpenStudy (solomonzelman):

Sinx/n is it equal to derivative of a 4?

OpenStudy (shamil98):

d/dx 4 = 0 ..

OpenStudy (solomonzelman):

so yeah.

OpenStudy (shamil98):

d/dx sin x = cos x

OpenStudy (shamil98):

if you're asking for the derivative of sinx/n then you would use the quotient rule

OpenStudy (solomonzelman):

what is it?

OpenStudy (shamil98):

f'g - g'f / g^2

OpenStudy (solomonzelman):

Woops, wrong thing

OpenStudy (shamil98):

let sin x = f(x) and n = g(x) so your derivative of sinx/n f'g-g'f / g^2 = cosx(n) - 1(sinx) / n^2

OpenStudy (solomonzelman):

I think I got til now.....

OpenStudy (shamil98):

f(x) = sin x f'(x) = cos x g(x) = n g'(x) = 1